OpenStudy (anonymous):

s=200, r=4000. If s=r*Theta, find theta in radians and degrees. If someone can check my answers That'd be great

OpenStudy (anonymous):

my answers are radians: pi/3600 degrees: 3'

OpenStudy (anonymous):

how did you solve it?

OpenStudy (anonymous):

in the formula \(s=r \theta\), the \(\theta\) is already in radians

OpenStudy (anonymous):

did you use the given formula?

OpenStudy (anonymous):

if we will plug-in the given values...\[200=4000\times \theta\]divide by 4000\[\theta=\frac{200}{4000}=0.05~rad\]

OpenStudy (anonymous):

to convert to degrees, there are \(180^\circ=\pi~rad\) so...\[0.05~\bcancel{rad}\times \frac{180^\circ}{\pi~\bcancel{rad}}=9^\circ\]